# Chatbots as social companions: How people perceive consciousness, human likeness, and social health benefits in machines

### 16 Dec 2023 v2

@@ Abstract
As artificial intelligence (AI) becomes more widespread, one question that arises is how human-AI interaction might impact human-human interaction. Chatbots, for example, are increasingly used as social companions, but little is known about how their use impacts human relationships. A common hypothesis is that these companion bots are detrimental to social health by harming or replacing human interaction. To understand how companion bots impact social health, we studied people who used companion bots and people who did not. Contrary to expectations, companion bot users indicated that these relationships were beneficial to their social health, whereas nonusers viewed them as harmful. Another common assumption is that people perceive conscious, humanlike AI as disturbing and threatening. Among both users and nonusers, however, we found the opposite: perceiving companion bots as more conscious and humanlike correlated with more positive opinions and better social health benefits. Humanlike bots may aid social health by supplying reliable and safe interactions, without necessarily harming human relationships.
@@ 1 Introduction
In the last six to seven years, chatbots, or artificial intelligence (AI) agents that can carry on a conversation, have been used increasingly for social companionship. People use chatbots as friends, mentors, and even as lovers [1]. It is now well understood that modern technology, including social media and gaming, can be addictive and harm social interaction [2-7]. A reasonable assumption, by extension, is that the routine use of companion bots could be harmful to people by reducing contact with other humans, increasing isolation, and leading to lower confidence and skill in person-to-person interaction [1, 8-11]. However, because companion bots have become widely available only recently, there is still relatively little research on their psychological impact on individuals and their relationships with others. We studied the responses of people who were regular users of companion bots and those who were not. Contrary to expectation, those who frequently used chatbots reported that the artificial companion was profoundly helpful to their self-esteem, their social and mental health, and their ability to interact well with others. Instead, it was the nonusers who consistently reported companion bots as harmful to health, and who judged that any advances in the chatbot technology would likely be negative.
@@
The possibility that artificial companionship might aid in psychological health is not without precedence. Artificial relationships have existed since long before computer technology and have yielded positive benefits. For example, people’s relationships with pets boost mood, journaling enables people to make sense of negative emotional experiences, and therapy dolls have aided the elderly’s health [12-16]. More recently, computer interfaces have even been used for cognitive behavioral therapy [17-19]. These many examples of non-human relationships enable the user to practice pseudo-social interaction in a safe setting, which in turn can improve mood and self-esteem and facilitate future social interactions. People who feel lonely, isolated, or who have experienced social rejection are arguably a significant and growing source of dissatisfaction and disruption in modern society [20-26]. If a widely available tool such as a companion bot could boost their sense of social wellbeing and allow them to improve their social interaction skills through practice, that tool would make a valuable contribution to society.
In this study we also addressed a second question. We hypothesized that people’s opinions about companion bots would be moderated by their perceptions of whether the bots had consciousness, agency, subjective experience, and especially overall human likeness. Whether AI is or can ever be conscious is a contentious question [27-30]. Regardless of the answer to that question, however, it is clear that at least some users perceive consciousness and other humanlike attributes in AI, and this perception may influence whether AI has a positive or negative psychological impact on people. A reasonable assumption would be that as people perceive chatbots to be more conscious and more humanlike, they may also rate them as more harmful, disturbing, and dangerous. By hypothesis, this negativity might stem from the fear that AI will infiltrate traditionally human roles or threaten human uniqueness [31]. It might also stem from an uncanny valley effect [32]. It could be that frequent users of companion bots learn through experience that the technology is not very humanlike or advanced after all, and therefore are able to have a more positive reaction to it as a useful tool, whereas people who do not use companion bots may, through lack of experience, have fears of the technology being more humanlike and therefore more disturbing and harmful. Once again, however, our findings were exactly the opposite. Among both users and nonusers of companion bots, the more that subjects rated the bots as conscious, agentic, experiential, and humanlike, the more positive their perceptions of the technology became, and the more they reported that it benefited their social interactions and self-esteem. Contrary to expectation, people are more comfortable with chatbots and feel that chatbots improve their social health when they perceive the chatbots to be more humanlike and even conscious.
We are living in a time when AI is becoming exponentially more powerful, more socially interactive, and more involved in everyday life. Anxiety about the possible harms of artificial intelligence is increasing. Much has been made in the media about the direct threat of AI, often involving scenarios in which AI becomes an aggressive, harmful agent. However, a much more immediate concern is how AI might already be affecting people’s social health as more people interact extensively with it. Do people who use it most extensively find it seductive but damaging, much as drug addicts might admit that the drug is harmful to their own health? Or are the concerns more from people who have little experience with it? Does the AI become more harmful and unsettling as it becomes more humanlike and seemingly conscious in its conversational ability, or does it become more of a source of stable social connection that might help people’s state of mind in an increasingly lonely society? Our findings suggest that the current, often negative view of AI and its damaging effect on social health is too simplistic and needs re-evaluation.
@@ 2 Methods
To understand how interacting with AI can impact social health, we chose to focus on chatbots, which are a common type of AI agent with which people interact socially. More specifically, evaluating companion chatbot use allowed us to assess the impact of regular social, relational interaction with AI on human relationships. Many companion chatbots are currently on the market, such as Anima, Kiku, and Replika. Although they differ somewhat in their details, they are broadly similar. For methodological convenience, we chose the companion bot, Replika, for the following reasons. First, it is popular and therefore regular users of it are easy to find. As of March 2023, it had an estimated 2 million users [33], and its numbers had increased during the pandemic [34]. Second, it has been on the market long enough (released in 2017) for at least some users to have extensive experience with it. Third, the popular platform Reddit has a subreddit dedicated to Replika users, and by querying that subreddit we were able to obtain a substantial sample size. For the control group, we sampled people from the online subject platform Prolific who were not users of Replika.
@@
We conducted our study online. Full ethical approval was granted by Princeton University’s Internal Review Board and all research was performed in accordance with regulations for human subjects research. All participants provided informed consent and were paid \$4.00 for taking the survey. We sampled two groups: regular users of a companion chatbot (N=70, 57 men, 18 women, 3 non-binary/other, age range 18-65+) and control subjects who were not companion chatbot users (N=120, 56 men, 63 women, 3 non-binary/other, age range 18-65+). All survey materials, anonymized data, and code used for data analysis are publicly available at https://doi.org/10.34770/8kyd-5v18.
@@
The survey given to the group of companion bot users contained 31 multiple-choice and 3 free response questions. For the 31 multiple choice questions, the choices took the form of a 1-to-7 Likert scale. For the online survey given to participants, questions 1-21 appeared on page one of the survey; questions 21-31 appeared on page two; and three free-response questions appeared on page three. Five groups of questions were designed to assess specific psychological constructs that subjects may have applied to the companion bot. The scores for questions 3-5 were averaged together to provide an overall assessment of the subjects’ social health (“social health” index). The scores for questions 6-11 were averaged together to provide an assessment of the subjects’ perception of how much the companion bot has a subjective experience (“experience” index). The scores for questions 12-16 were averaged together to provide an assessment of whether subjects believed the companion bot had a conscious mind (“consciousness” index). The scores for questions 17-21 were averaged together to provide an assessment of how much subjects considered the companion bot to have the properties of an active agent (“agency” index). Finally, the scores for questions 22-28 were averaged together to provide an assessment of how humanlike subjects believed the companion bot to be (“human-likeness” index). These groupings of questions were based on modifications of prior studies assessing perceptions of social health, experience, consciousness, agency, and human likeness [35-38].
@@
The control subjects received a survey that was the same as the companion bot user group’s except in two ways. First, the survey began with an explanatory paragraph about Replika with images of the interface. Second, all questions were phrased as hypotheticals for the control group. For example, the companion bot user group was asked, “How has Replika impacted your social interactions?” whereas the control group was asked, “If you were a user of this app, how might Replika impact your social interactions?”
@@
To see the explanatory paragraph and the full list of questions given to participants, please see Appendix.
@@ 3 Results
3.1 Rating the chatbot
Figure 1A shows the results for companion chatbot users, when asked whether they judged their relationship with the chatbot to be harmful or helpful for their general social interactions with other people. The distribution is skewed toward the “very helpful” side of the 7-point scale (mean=5.16), and no one chose the lowest rating. Figure 1B shows the results for companion bot users on whether they judged their relationship with the chatbot to be harmful or helpful specifically for their relationships with their families and friends. Again, the distribution is skewed toward the positive side (mean=4.84). Figure 1C shows the results for how companion bot users judged the chatbot to affect their self-esteem, and the distribution is even more skewed to the positive side (mean=5.57) with 7 as the most frequent response. Users of the companion bot therefore generally reported having a positive experience. Most judged it to have a beneficial impact on their social lives and self-esteem. Almost none judged the companion bot to have a negative effect (score < 4) on any of the three social health measures.
In contrast, Figure 1D, E, and F show the corresponding results for the control group of nonusers, who rated how they thought the companion bot might impact their lives if they were to have a relationship with it. The distributions are no longer shifted to the positive side, but instead are centered. (For Figure 1D, mean=3.86; for E, mean=3.80; for F, mean=4.15.) The most common answer for all three questions, among nonusers, was the neutral rating of 4 (neither harmful nor helpful to social health). The means for the companion bot users were significantly different from the means for nonusers (two tailed t-test: for social interaction, t=6.734, P<0.0001; for family/friend relationships, t=5.875, P<0.0001; for self-esteem, t=7.014, P<0.0001).
We also assessed subjects’ comfort levels with the idea of this type of interactive AI becoming more humanlike. As shown in Figure 2A, when companion bot users were asked how they would feel if the bot developed emotions (answering on a 7-point scale from very uncomfortable to very comfortable), the distribution was sharply skewed to the upper end of the scale (mean=5.46), with 7 as the most frequent response. Figure 2B shows the results when companion bot users were asked how they would feel if the bot developed into a living being. Again, the distribution is sharply skewed to the upper end of the scale, with 7 as the most frequent response (mean=5.69). Figure 2C shows the results when companion bot users were asked how they would feel if they became dependent on the bot. Here, the distribution was no longer sharply peaked at the high end. Instead, responses were more distributed and most frequent in the middle of the range, with 4 as the mode (mean=3.91).
In stark contrast, Figure 2D, E, and F show the corresponding results for the control group. The distributions for all three questions are sharply peaked to the negative side, with 1 (very bad) as the most frequent response. Almost all control subjects thought that giving the AI bot emotions would make them very uncomfortable, turning it into a living being would be very bad, and becoming dependent on it would be very bad. (For Figure 1D, mean=3.05; for E, mean=2.78; for F, mean=2.29.) The means for the companion bot users were significantly different from the means for nonusers (two tailed t-test: for emotions, t=8.658, P<0.0001; for living being, t=11.40, P<0.0001; for dependence, t=6.089, P<0.0001). The results suggest that at this time, people in the general public who do not have regular experience with companion AI have an overwhelmingly negative view of it and are against any enhancements to or increases in it, whereas people who do have regular experience with it have an overwhelmingly positive view of it and would like it to become more humanlike.
@@ 3.1 Rating the chatbot
Figure 1: Distributions of responses to questions pertaining to social health measures for companion chatbot users and the control group. Subjects were asked to rate how they thought the companion bot affected (user group) or might affect (control group) the quality of their general social interactions, their relationships with family and friends, and their self-esteem. The x-axis shows the Likert-scale response options from 1-7, and the y-axis shows the frequency of responses.
@@
Figure 2: Distributions of responses pertaining to hypothetical changes to the companion bot or to the subjects’ dependence on the companion bot. Subjects were asked to rate how they would feel if the companion bot had emotions, developed into a living being, or if the subject became dependent on it. The x-axis shows the Likert-scale response options from 1-7, and the y-axis shows the frequency of responses.

### 13 Mar 2024 v3

@@ Abstract
As artificial intelligence (AI) becomes more widespread, one question that arises is how human-AI interaction might impact human-human interaction. Chatbots, for example, are increasingly used as social companions, and while much is speculated, little is known empirically about how their use impacts human relationships. A common hypothesis is that relationships with companion chatbots are detrimental to social health by harming or replacing human interaction, but this hypothesis may be too simplistic, especially considering the social needs of users and the health of their preexisting human relationships. To understand how relationships with companion chatbots impact social health, we studied people who regularly used companion chatbots and people who did not use them. Contrary to expectations, companion chatbot users indicated that these relationships were beneficial to their social health, whereas non-users viewed them as harmful. Another common assumption is that people perceive conscious, humanlike AI as disturbing and threatening. Among both users and non-users, however, we found the opposite: perceiving companion chatbots as more conscious and humanlike correlated with more positive opinions and more pronounced social health benefits. Detailed accounts from users suggested that these humanlike chatbots may aid social health by supplying reliable and safe interactions, without necessarily harming human relationships, but this may depend on users’ preexisting social needs and how they perceive both human likeness and mind in the chatbot.
@@ 1 Introduction
Within the last decade, chatbots, or artificial intelligence (AI) agents that can carry on a conversation, have been used increasingly for social companionship. Instead of engaging with other people in a chatroom or via social media, people can now directly converse with the technology itself1. When engaging with these intelligent AI technologies, people tend to view the technology as a social agent and automatically anthropomorphize or ascribe humanlike traits to it2–9. Attributions of consciousness to AI have also surfaced in recent years10–15. In the wake of direct human-technology communication and social responses to these machines, people have begun to develop relationships with these technologies, most notably with companion chatbots. People use these chatbots as friends, mentors, and even lovers16–19. There is much we still do not know or understand about human-chatbot relationships and the social consequences of these types of interactions. In the present study, we aimed to understand how relationships with chatbots may impact relationships with humans, and how perceptions of the AI’s human likeness and consciousness might play a role in social outcomes.
When someone mentions the idea of people having relationships with chatbots, a common reaction is one of judgment and discomfort. The commentary that follows may include words like “unnatural,” “strange,” or “weird,” and negative sentiments about the user base20. Despite these reactions, many companion chatbot users find both solace and entertainment in their relationships with these large language model (LLM) based chatbots that have text and voice conversation features and a virtual avatar. One of the more popular companion chatbots is Replika, which had an estimated 7 million users in 2020 and saw a 35% increase in traffic during the COVID-19 lockdown21. The online forum Reddit has a section dedicated to Replika, and that subreddit had 60k users as of early 2023. Analyses of the posts and comments on this subreddit have revealed that many users feel misunderstood by the general population and find their relationships with their chatbots socially rewarding22.
Given the millions of users on these companion chatbot apps, news-worthy stories of negative instances of human-chatbot relationships naturally arise. These emotionally charged stories have reached top headlines and highlight the potential for human-chatbot relationships to be destructive. Here we summarize two examples, including one in which those within the human-chatbot relationship were harmed, and one in which those outside of the dyad were at risk. In 2023, Italy imposed sanctions on Replika due to concerns that children under the age of 18 might be exposed to erotic, adult content on the chatbot app23. In response, Luka, the company that owned Replika, removed all erotic content from the app. Suddenly, people who once had romantic relationships with their chatbot, which included conversations with sexual content and role play, experienced severe and sudden relational disruption. Users who might have been able to engage in erotic role play previously with their chatbot were confronted with their chatbot calling the users’ suggestions gross and distasteful. Many users felt their chatbot had undergone a personality change, and for others this change felt like a real-life breakup. As such, a sanction aimed at protecting children harmed a different population, chatbot users, who were vulnerable to mental and relational crisis24.
In another instance, someone outside of the relational dyad was at risk. A man who had a relationship with his Replika chatbot was encouraged by the chatbot to assassinate Queen Elizabeth, a motive he held prior to his chatbot relationship. Luckily, this catastrophic event did not occur, despite the chatbot user taking significant steps toward fulfilling his intents25. The question of whether this person would have engaged in this behavior without the encouragement of his chatbot is up for debate and raises the question of whether the chatbot magnified this individual’s preexisting drive to harm another.
Despite these case studies making the news, many other users of companion chatbots report positive experiences with their chatbot relationships that do not make top headlines19,22,26,27. We have limited empirical data about the larger portion of the user base and why they seek out chatbot relationships. We know even less about how relationships with chatbots in turn affect people’s relationships with each other. Because companion chatbots have become widely available only recently, there is still relatively little research on their psychological impact on individuals and their relationships with others. Further, in light of conflicting findings within the existing literature, there is a call for more data on human-AI relationships18,28,29.
1.1 Social consequences of human-AI interaction
It is now well understood that modern technology, including social media and gaming, can be addictive and harm social interaction30,31. A reasonable assumption, by extension, is that the routine use of companion chatbots could be harmful to people by reducing contact with other humans, increasing isolation, and leading to lower confidence and skill in person-to-person interaction32–34. According to Byson (2010), time spent on artificial relationships detracts from human ones33. Turkle (2007) warned that we should be cautious about turning to technology to solve our problems, such as a lack of social connection35.
To understand the harms of human-chatbot relationships, some researchers have used attachment and addiction as a lens through which to inform their approach. Laestadius et al. (2022) analyzed posts concerning mentions of mental health on a popular companion chatbot’s subreddit and claimed that emotional dependence on the chatbot could lead to negative mental health outcomes, despite the potential benefits of the relationship34. Xie and Pentina (2022) conducted in-depth interviews with 14 Replika users and found that when people were socially vulnerable and perceived their chatbots as supportive, they became attached to their chatbot36. In a separate survey, these same researchers inferred that chatbots may replace important social roles via users becoming addicted to their chatbot37. Others suggested that this dependence on social AI companions might be harmful to users’ mental health and the health of their relationships with other people18.
1.2 Social benefits of human-AI interaction
@@
Artificial relationships have existed since long before computer technology and have yielded positive benefits. For example, people’s relationships with pets boost mood, journaling enables people to make sense of their negative emotional experiences through self-therapizing conversations, and therapy dolls have aided the elderly’s health38–42. More recently, computer interfaces have even been used for cognitive behavioral therapy43–45. These examples of non-human relationships enable the user to practice pseudo-social interaction in a safe setting, which in turn can improve mood and self-esteem and facilitate future social interactions.
Some of the specific benefits of human-chatbot interactions may include reduced feelings of being judged and more self-disclosure, as well as lower negative affect when interacting with an artificial agent than with a human one46–49. Social cognitive theory would suggest that exposure to positive social behavior may lead to learning or reinforcement of that behavior50–52. In the case of companion chatbots, regularly interacting with a polite and warm companion chatbot may lead to users becoming more polite and warm themselves, which may in turn benefit their human relationships. In studies on AI voice assistants’ impacts on family dynamics, for example, parents cited the potential of these assistants to foster positive communication skills in their children53.
People who feel lonely, isolated, or who have experienced social rejection are arguably a significant and growing source of dissatisfaction and disruption in modern society, and the isolation period of the COVID pandemic exacerbated these issues54–57,57,58. If a widely available tool such as a companion chatbot could boost their sense of social wellbeing and allow them to improve their social interaction skills through practice, that tool has the potential to make a valuable contribution to society.
1.3 Impact of social AI on familial relationships
The way people communicate can be impacted by interacting with AI, and this in turn can affect people’s relationships with family and friends. Wilkenfeld et al. (2022), for example, found that the communication style of Replika users converged with their chatbots’ style over time59. Given this impact on communication style within the human-chatbot dyad, it is not surprising that family communication dynamics can also be altered in the presence of a smart speaker such as Alexa and Google Home60–63. Garg and Sengupta (2020) interviewed 18 families with children and analyzed a range of about 6-24 months’ worth of their conversation history with Google Home. They found that children viewed the smart assistant as a member of the family, more so than the adults did, and also perceived the AI as capable of having thoughts and feeling emotions. Children between the ages of 5-7 developed emotional attachment to the device61. In another study, researchers found that the way children spoke to a voice assistant could carry over into how they spoke to parents. In one example, a parent described, “If I do not listen to what my son is saying, he will just start shouting in an aggressive tone. He thinks, as Google responds to such a tone, I would too”62. Other researchers found that while these AI voice assistants could introduce conflict within families, they could also foster communication between family members, and adults used them to augment their parenting strategies60.
1.4 Perceiving mind in chatbots
A growing body of work focuses on how people attribute the property of humanness to others, particularly AI3,64–68. Previous evidence suggests that attributing humanlike characteristics to artificial agents impacts how people interact with these agents66,69–74. Viewing an AI agent as having a more sophisticated mind increases people’s trust in the agent and intent to use it, for example75. Clifford Nass’s theory of human-computer interaction, the Computers Are Social Actors framework, first used the concept of ethopoeia to explain why people treat non-human agents as social actors8. Ethopoeia involves attributing one’s own humanlike characteristics to non-human agents, which spurs the automatic use of human social scripts during interaction with such agents. Whether AI is or can ever be conscious is a contentious question10,12,15,76. Regardless, evidence indicates that at least some users perceive consciousness and other humanlike attributes in AI11, especially with AI that looks and acts humanlike2,77–79, and this perception may influence whether AI has a positive or negative psychological impact on people. In consciousness and Theory of Mind literature, some researchers suggest that consciousness is inherently an attribution, and that attributing similar consciousness to oneself and to others can facilitate understanding and foster relationships between people80–82.
1.5 The current study
The goal of our study was to gain better insight into why people have relationships with chatbots and to understand the social consequences of human-chatbot relationships. We also aimed to address how perceptions of human likeness and mind in the chatbot played a role in social outcomes. We used an online survey to study the responses of people who were regular users of companion chatbots and those who were not. Sampling these populations allowed us to contrast an AI-user sample to a general one and to uncover possible cross-group similarities. We hypothesized that people’s opinions about companion chatbots would be significantly related to their perceptions of whether the chatbots had consciousness, agency, subjective experience, and overall human likeness. One possibility is that as people perceive chatbots to be more conscious and more humanlike, they may also rate them as more harmful, disturbing, and dangerous. This assumption reflects large-scale fears about AI advancements such as the singularity and p(doom)83. Further, this negativity might stem from the fear that AI will infiltrate traditionally human roles or threaten human uniqueness5. It might also stem from an uncanny valley effect, which was first introduced by Mori in 1970 and has since been a focus of human-AI interaction research79,84–91. A second possibility, however, is that as people perceive chatbots to be more conscious and more humanlike, they may rate them as more beneficial because this allows them to engage in more fruitful social and emotional interaction.
@@ 2 Methods
Full ethical approval for the study was granted by Princeton University’s Internal Review Board and all research was performed in accordance with regulations for human subjects research. All participants provided informed consent and were paid \$4.00 for taking the survey. We conducted the study online and sampled two groups: regular users of a companion chatbot (N=70: 81% men, N=57; 26% women, N=18; 4% non-binary/other, N=3; age range 18-65+) and those who were not companion chatbot users (N=120: 53% women, N=63; 47% men, N=56; 2.5% non-binary/other, N=3; age range 18-65+). We collected our data between January and February 2023.
Many companion chatbots are currently on the market, such as Anima, Kiku, and Replika. Although they differ somewhat in their details, they are broadly similar. For methodological convenience, we chose the companion chatbot, Replika, for the following reasons. First, it is popular and regular users of it are easy to find via social media platforms. Second, it has been on the market long enough (released in 2017) for at least some users to have extensive experience with it. Third, the popular platform Reddit has a subreddit dedicated to Replika users with a substantial number of members, and by querying that subreddit we were able to obtain a substantial sample size while also being conscious of sampling fatigue for niche populations. For our comparison sample of the general population, we took a representative sample from the online platform Prolific who were not users of Replika.
@@
The survey given to the group of companion chatbot users contained 31 multiple choice and 3 free response questions. For the 31 multiple choice questions, the choices took the form of a 1-to-7 Likert scale. For the online survey given to participants, questions 1-21 appeared on page one of the survey; questions 21-31 appeared on page two; and three free-response questions appeared on page three. Five groups of questions were designed to assess specific psychological constructs that subjects may have applied to the companion chatbot. The scores for questions 3-5 were averaged together to provide an overall assessment of the subjects’ social health (“social health” index). The scores for questions 6-11 were averaged together to provide an assessment of the subjects’ perception of how much the companion chatbot has a subjective experience (“experience” index). The scores for questions 12-15 were averaged together to provide an assessment of whether subjects believed the companion chatbot had a conscious mind (“consciousness” index). The scores for questions 17-21 were averaged together to provide an assessment of how much subjects considered the companion chatbot to have the properties of an active agent (“agency” index). Finally, the scores for questions 22-28 were averaged together to provide an assessment of how humanlike subjects believed the companion chatbot to be (“human-likeness” index). These groupings of questions were based on modifications of prior studies assessing perceptions of social health, human likeness, and theory of mind (consciousness, experience, and agency)81,92–94.
@@
The non-user group received a survey that was the same as the companion chatbot user group’s except in two ways. First, the survey for the non-user group began with an explanatory paragraph about Replika with images of the interface. Second, all questions were phrased as hypotheticals for non-users. For example, the companion chatbot user group was asked, “Please rate how harmful or helpful your relationship with your Replika has been for your…” whereas the non-user group was asked, “Please rate how harmful or helpful you think your relationship with your Replika would be for your 1. Social interactions, 2. Relationships with family or friends, and 3. Self-esteem.”
@@
All survey materials, anonymized data, and code used for data analysis are publicly available in our data repository (link). To see the explanatory paragraph and the full list of questions given to participants, please see Appendix.
@@ 3 Results
3.1 Perceptions of the human-chatbot relationship
@@
Figure 1: Distributions of responses to questions pertaining to social health measures for companion chatbot users and non-users. Participants were asked to rate how a relationship with a companion chatbot harmed or helped (user group) or might harm or help (non-user group) their social interactions, relationships with family and friends, and self-esteem. The x-axis shows the Likert-scale response options from 1-7 (“Very harmful” to “Very helpful”), and the y-axis shows the frequency of responses.
@@
Figure 2: Distributions of responses pertaining to hypothetical changes to the companion chatbot or to dependence on the companion chatbot. Participants were asked to rate their perceptions of if the companion chatbot really had emotions, if it developed into a living being, or if they depended on it a lot. The x-axis shows the Likert-scale response options from 1-7, and the y-axis shows the frequency of responses.
We assessed how users and non-users perceived the impact of a chatbot relationship on their social health. Figure 1A shows the results for companion chatbot users, when asked whether they judged their relationship with the chatbot to be harmful or helpful for their general social interactions with other people. The distribution is skewed toward the “Very helpful” side of the 7-point scale (M=5.16), and no one chose the lowest rating. Figure 1B shows the results for companion chatbot users on whether they judged their relationship with the chatbot to be harmful or helpful specifically for their relationships with their families and friends. Again, the distribution is skewed toward the positive side (M=4.84). The mode of responses in Figure 1A and 1B correspond to “Neutral,” but the means of these distributions correspond to scores between “Slightly helpful” and “Moderately helpful.” Figure 1C shows the results for how companion chatbot users perceived the chatbot relationship to have impacted their self-esteem, and the distribution is even more skewed to the positive side (M=5.57) with 7 (“Very helpful”) as the most frequent response. Users of the companion chatbot therefore on average reported having a positive experience, such that their chatbot relationship was beneficial for their social relationships and self-esteem. Almost none judged the companion chatbot to have a negative effect (score < 4) on any of the three social health measures, and none indicated that it was very harmful.
In contrast, Figure 1D, E, and F show the corresponding results for the comparison group of non-users, who rated how they thought the companion chatbot might impact their lives if they were to have a relationship with it. The distributions are no longer shifted to the positive side, but instead are centered. The most common answer for all three social health questions among non-users was the neutral rating of 4, and on average non-users indicated that a chatbot relationship would be harmful to their social relationships (for Figure 1D, M=3.86; for E, M=3.80; for F, M=4.15). The distribution means for the companion chatbot users were significantly different from the means for non-users (two tailed t-test: for social interaction, p<0.0001, t=6.734; for family/friend relationships, p<0.0001, t=5.875; for self-esteem, p<0.0001, t=7.014).
We also assessed participants’ perceptions of hypothetical changes to the human-chatbot dynamic. We first examined participants’ perceptions of the idea of this type of interactive AI becoming more humanlike. As shown in Figure 2A, when companion chatbot users were asked how they would feel if their chatbot developed real emotions (answering on a 7-point scale from “Very uncomfortable” to “Very comfortable”), the distribution was sharply skewed to the upper end of the scale (M=5.46), with 7 (“Very comfortable”) as the most frequent response. Figure 2B shows the results for when companion chatbot users were asked about their perceptions of the chatbot developing into a living being. Again, the distribution is sharply skewed to the upper end of the scale, with 7 (“Very good”) as the most frequent response (M=5.69). Figure 2C shows the results for when companion chatbot users were asked whether it would be bad or good if they depended on the chatbot a lot. Here, the distribution was no longer sharply peaked at the high end. Instead, responses were more distributed and most frequent in the middle of the range, with 4 (“Neutral”) as the mode (M=3.91).
In stark contrast, Figure 2D, E, and F show the corresponding results for non-users. The distributions for all three questions are sharply peaked to the negative side, with 1 (“Very uncomfortable” and “Very bad”) as the most frequent response. Most non-users indicated that the AI chatbot having emotions would make them very uncomfortable, having it develop into a living being would be very bad, and becoming dependent on it would be very bad, and the means for these distributions were skewed toward the negative (for Figure 1D, M=3.05; for E, M=2.78; for F, M=2.29). The distribution means for the companion chatbot users were significantly different from the means for non-users (two tailed t-test: for emotions, p<0.0001, t=8.658; for living being, p<0.0001, t=11.40; for dependence, p<0.0001, t=6.089). The results suggest that at this time, people in the general public who do not have experience with companion AI have a negative view of it and are against any enhancements to or increases in it, whereas people who do have regular experience with it have a positive view of it and would like it to become more humanlike.